1. You have a 0.02 M stock solution of Tris base. How would you make 500 ml of 0.3 mM Tris base ?


1
Expert's answer
2021-10-05T01:36:10-0400

CM = n / V

n1 = n2

n = CM x V

n2 = 0.3/1000 x 500/1000 = 0.00015 mol

n1 = n2 = 0.00015 mol

V1 = n1/CM1 = 0.00015/0.02 = 0.0075 L = 7.5 ml

In order to make 500 ml of 0.3 mM Tris base, 7.5 ml of a 0.02 M stock solution must be diluted by distilled water to the volume of 500 ml.